### v1

# FLawN-T5: An Empirical Examination of Effective Instruction-Tuning Data Mixtures for Legal Reasoning

Abstract
Instruction tuning is an important step in making language models useful for direct user interaction. However, many legal tasks remain out of reach for most open LLMs and there do not yet exist any large scale instruction datasets for the domain. This critically limits research in this application area. In this work, we curate LawInstruct, a large legal instruction dataset, covering 17 jurisdictions, 24 languages and a total of 12M examples. We present evidence that domain-specific pretraining and instruction tuning improve performance on LegalBench, including improving Flan-T5 XL by 8 points or 16% over the baseline. However, the effect does not generalize across all tasks, training regimes, model sizes, and other factors. LawInstruct is a resource for accelerating the development of models with stronger information processing and decision making capabilities in the legal domain.
1 Introduction
In recent years, Large Language Models (LLMs) advanced significantly, evident in their performance gains across numerous benchmarks, including SuperGLUE (Wang et al., 2019), MMLU (Hendrycks et al., 2021b), and various human examinations (OpenAI, 2023), such as the U.S. bar exams for law practice admission (Katz et al., 2023). However, the interplay between domain-specific training and within-domain evaluation is poorly understood. This work examines how training on domain-specific legal corpora affects performance on the widest set of legal-domain evaluation benchmarks known to the authors. We thus conduct a study of the ability of models to answer questions, classify, make judgments, extract information, and otherwise perform decision making or higher-order cognitive tasks (i.e., to “reason”) within a limited domain, as opposed to broad-domain benchmarking. We present evidence that domain-specific pretraining and instruction tuning improve performance—but the effect does not generalize across all tasks, training regimes, model sizes, and other factors.
Although large closed models also still hallucinate heavily on legal texts (Dahl et al., 2024), they achieve much better performance on LegalBench than smaller open models (e.g., 77.3 for GPT-4 vs. 60.1 for Flan-T5 XXL, the state-of-the-art open model). In the legal domain it is often crucial for reasons of trust and data protection not to use public models, so many firms need on-premise deployments. Therefore models like Claude or GPT-4 cannot be used, stressing the need for open models. In this study, we explore the potential of enhancing model performance through in-domain instruction tuning and continued pretraining on Flan-T5, the current state-of-the-art open model on LegalBench in both the 3B and 11B range.
Figure 1: We continue pretraining on MultiLegalPile, instruction tune on LawInstruct and evaluate on LegalBench and MMLU.
To study this, we use the MultiLegalPile (Niklaus et al., 2023b), a 689GB multilingual legal corpus, for continued pretraining. Because no instruction dataset for legal reasoning is available, we introduce LawInstruct, spanning 24 languages in 17 jurisdictions on four continents. It contains 12M training examples for QA, entailment, summarization, and information extraction tasks in the legal domain, each presented as a bespoke instruction with corresponding output. With this large instruction dataset in hand, we fine-tune models and then perform quantitative analyses of their outputs on the LegalBench (Guha et al., 2023) and MMLU (Hendrycks et al., 2021a) benchmark suites. Instruction tuning Flan-T5 models on LawInstruct, we achieve a balanced accuracy of 58.1 on LegalBench for the XL size, improving by 8 points or 16% over the baseline. The Small model even improves by 9.6 points or 38.1% and by 14 points or 55.4% when we also continue pretraining it.
The contributions of this paper are four-fold: First, we curate the first legal instruction dataset by unifying and writing instructions for 58 high-quality annotated datasets covering diverse legal tasks. Second, we continue pretraining and instruction tune T5, mT5, and Flan-T5 models and achieve new state-of-the-art on LegalBench in all tested parameter ranges. Third, we perform a wide range of ablations across different dataset configurations providing new insights for domain adaptation. Finally, we publicly release the permissively-licensed portion of the curated dataset on the Hugging Face Hub [^1] https://huggingface.co/lawinstruct/lawinstruct and release the code used to create the dataset [^2] https://github.com/JoelNiklaus/LawInstruct/ including pointers on how to access the portions of the data that require special agreements.
2 Experimental Setup
In this section, we describe the experimental setup we used to test the effect of pretraining and instruction tuning on in-domain legal data. We consider T5 v1.1+LM adaptation (Raffel et al., 2020; Lester et al., 2021), Flan-T5 (Chung et al., 2022) and mT5 (Xue et al., 2021) models in the sizes Small, Base, XL and XXL, allowing us to study effects over different model scales. We selected the T5 family of models over other models for three reasons: 1) Flan-T5 XL and XXL perform best in their parameter range on LegalBench, 2) T5 and mT5 allow us to measure the effect of multilinguality in a controlled setting, and 3) the T5 model family contains models from 60M parameters (Small) to 11B (XXL) allowing us to study scaling behaviour also at smaller scales.
We used random seed 42 throughout. Our experiments were performed with T5X  [^3] https://github.com/google-research/t5x on TPUv4 pods using between 2 and 512 cores. We present the mean across tasks per LegalBench category and for LegalBench overall by aggregating over the categories.
2.1 Continued Pretraining
We continue pretraining on the MultiLegalPile (Niklaus et al., 2023b), a 689GB corpus in 24 languages from 17 jurisdictions. The MultiLegalPile corpus, which includes diverse legal data sources with varying licenses, allows for pretraining NLP models under fair use, with more permissive licenses for the Eurlex Resources and Legal mC4 subsets. It consists of four large subsets: a) Native Multi Legal Pile (112 GB), b) Eurlex Resources (179 GB), c) Legal mC4 (106 GB), and d) Pile of Law (292 GB). For our experiments on the mT5 models, we use the entire corpus, and for the experiments on the T5 and Flan-T5 models we use only the texts in the English language.
We continued pretraining (a.k.a. domain adaptation of) the models with 512 tokens in both inputs and targets on the MultiLegalPile (Niklaus et al., 2023b) whereas the original models were pretrained on C4 (Raffel et al., 2020). We used the UL2 mixture (Tay et al., 2022) due to its promise to enable improved training efficiency with its mixture of denoisers tasks. In initial experiments we used batch size 1024 and warmed up the learning rate linearly for the first 10K steps from 2.5e-3 to 5e-3 and finally decayed down to 1.5e-3. However, we noticed training instabilities for the XXL models. We switched to a constant learning rate of 1e-3 and ran a sweep over batch sizes 64, 128, 256, 512, 1024. The XXL model only trained stably with batch size 128.
2.2 Instruction Tuning
[…]
Instruction: Answer the following legal question. Cite relevant evidence when possible. Prompt: Question: So say a bank robber walks into a bank and hands the teller a note saying simply "please give me \$1,000" (or maybe even just verbally saying "I need \$1,000 please") and if the teller accepts, which they probably would because they’re supposed to comply with the robber, he takes the money, otherwise if they refuse, he leaves. How is this different than say people asking for charity donations in the street? He just politely requested money without making any threats. Answer: Answer: As a concrete example, consider Missouri v. Coleman, where Coleman handed a teller a plastic bag and said "I need you to do me a favor. Put the money in this bag", and later as the manager approached, said "Ma’am, stop where you are and don’t move any farther". Coleman was convicted of second degree robbery: but the appeals court found that he had not acted forcibly, so his conviction was overturned. Instead, the court entered a conviction for the lesser offense of stealing which is when one appropriates property or services of another with the purpose to deprive him or her thereof, either without his or her consent or by means of deceit or coercion. The act may also be termed "theft", as in Washington state. The thief is acting deceptively and thereby gaining control over property.
Figure 2: Instruction template in LawInstruct for StackExchangeQuestionsLegal populated with instruction, prompt and answer. Models are trained to generate the answer conditioned on the instruction and prompt.
Figure 3: Jurisdiction and task type by examples.
(a) Instructions (capped at 100)
(b) Prompts (capped at 15K)
(c) Answers (capped at 1K)
Figure 4: Mean length distributions for instructions, prompts and answers.
We finetuned the models with 2048 input and 512 target tokens. We ran a hyperparameter sweep for the XL model over the learning rate (5e-5, 1e-5, 5e-4, 1e-4, 5e-4) and dropout (0, 0.05, 0.1, 0.15, 0.2, 0.25), with learning rate 5e-4 and dropout 0.15 achieving the best validation loss. Unless specified otherwise, we trained the models for 2K steps with batch size 64. In addition to LawInstruct, we used an updated Flan mixture (Chung et al., 2022). We built the input by concatenating the prompt with two new lines, the instruction and two additional new lines. Per LawInstruct config, we used the first 16 examples for validation and the remaining ones for training. We selected the model with the best LawInstruct validation loss. While in-context learning has achieved strong results in many tasks (Brown et al., 2020), further finetuning language models for specific tasks may still be necessary for better results (Mosbach et al., 2023).
2.3 Evaluation
We evaluate our models on LegalBench and MMLU to test in-domain and generalization performance, respectively. LegalBench (Guha et al., 2023) consists of 162 tasks evaluating different aspects of legal classification and reasoning. Each task is assigned to one of six categories, depending on the broader type of legal reasoning implicated. LegalBench tasks are sourced from both previously constructed datasets and novel tasks collected from different members of the legal community (e.g., lawyers, legal impact organizations, legal academics). As such, LegalBench is thought to capture tasks of interest and practical applicability. LegalBench tasks span a wide range of legal subject areas (e.g., contracts, civil procedure, tax, etc.) and text-types (natural language, contractual terms, judicial opinions, etc.). The majority of tasks are either classification or extraction tasks, thus enabling automated evaluation. Massively Multilingual Language Understanding (MMLU) benchmarks models factual knowledge (Hendrycks et al., 2021a). MMLU contains multiple-choice questions on 57 subjects, including three related to law: jurisprudence, international law, and professional law. While there exist multilingual benchmarks like LEXTREME (Niklaus et al., 2023a), they are still very challenging for generative models not specifically finetuned per task. Therefore, we focus on LegalBench and MMLU both in the English language.
For evaluation, we set temperature to 0 in line with accepted practice for LegalBench evaluation (Guha et al., 2023) that focuses on the highest-likelihood token sequence with minimal variance. We removed the following prefixes before scoring: “label”, “target”, “option”, “answer”, “a:”. We did not evaluate Rule QA because it necessitated manual evaluation. We show paper baseline results compared with our runs in Appendix D Table 3. Our XL model is quite close to the XL model in the LegalBench paper, but there are significant differences for the XXL model. We provide a more detailed analysis of possible causes in Section B.1. Unless specifically mentioned, we compare to our baselines results. We hold out LegalBench tasks overlapping with LawInstruct tasks unless specified otherwise (see Section B.2 for details).
3 Results
In this section, we discuss the main results achieved by instruction tuning and continued pretraining.
(a) LegalBench
(b) MMLU
Figure 5: Performance progression on LegalBench and MMLU from baseline to instruction finetuning and domain adaptation + instruction finetuning.
Table 1: Progression of performance from baseline to instruction tuning (IFT) and instruction tuning together with continued pretraining (PRE-IFT+IFT).
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench Improvement
Small Baseline 0.3 ± 0.7 30.4 ± 20.3 39.8 ± 20.8 28.2 ± 21.6 27.7 ± 21.9 25.3 ± 14.8 -
Small IFT 25.0 ± 22.0 38.1 ± 25.4 43.0 ± 17.1 36.1 ± 26.5 32.6 ± 24.2 34.9 ± 6.7 9.6 (38.1%)
Small PRE-IFT+IFT 51.6 ± 2.7 37.7 ± 25.2 39.8 ± 18.4 33.7 ± 23.3 33.8 ± 22.4 39.3 ± 7.4 14.0 (55.4%)
Base Baseline 44.7 ± 12.4 18.0 ± 23.6 20.9 ± 24.8 28.9 ± 21.2 37.0 ± 21.3 29.9 ± 11.1 -
Base IFT 50.3 ± 2.4 38.8 ± 25.9 40.5 ± 15.7 49.5 ± 19.1 45.2 ± 22.0 44.9 ± 5.2 15.0 (50.2%)
Base PRE-IFT+IFT 51.6 ± 4.8 38.2 ± 25.5 44.0 ± 13.4 45.4 ± 16.5 44.1 ± 19.0 44.6 ± 4.8 14.8 (49.5%)
XL Baseline 53.5 ± 6.0 32.1 ± 24.6 46.8 ± 15.6 58.7 ± 21.3 59.6 ± 25.6 50.1 ± 11.3 -
XL IFT 65.7 ± 15.2 45.1 ± 30.3 49.5 ± 14.2 61.7 ± 17.1 68.6 ± 24.1 58.1 ± 10.3 8.0 (16.0%)
XL PRE-IFT+IFT 60.3 ± 10.6 44.3 ± 29.7 50.5 ± 15.4 57.3 ± 15.9 67.3 ± 23.1 55.9 ± 8.9 5.8 (11.6%)
XXL Baseline 36.1 ± 21.5 18.8 ± 24.6 25.2 ± 26.0 35.1 ± 22.2 41.1 ± 18.4 31.3 ± 9.1 -
XXL IFT 55.2 ± 23.7 46.3 ± 31.6 56.2 ± 18.3 66.3 ± 19.7 73.8 ± 24.4 59.6 ± 10.6 28.3 (90.5%)
XXL PRE-IFT+IFT 52.2 ± 14.7 47.4 ± 30.8 59.2 ± 18.3 66.6 ± 18.5 70.0 ± 24.1 59.1 ± 9.5 27.8 (89.0%)
(a) LegalBench
(b) MMLU
Figure 6: Difference to the baseline for the XL model across tasks on LegalBench and MMLU. For LegalBench, we excluded tasks with a difference between -10 and 10 for clarity.
Figure 5 and Table 1 show the performance progression from the baseline over instruction tuning to domain adaptation + instruction tuning on LegalBench and MMLU. Instruction tuning leads to a large performance increase for all model sizes (38.1% for Small, 50.2% for Base, 16% for XL, and 90.5% for XXL). Domain adaptation + instruction tuning only improves further for the Small model size (55.4% vs. 38.1%). It seems like larger models benefit less from in-domain pretraining than smaller models, possibly because they can “remember” more from the pretraining phase due to increased capacity. Alternatively, a reason for non-consistent improvements of domain adaptation could be the switch from the UL2 tasks in continued pretraining to standard next-token prediction in instruction tuning. Finally, we conjecture that the switch from input length 512 tokens in continued pretraining to 2048 tokens in instruction tuning could have led lower performance for domain-adapted models.
(a) LegalBench
(b) MMLU
Figure 7: Difference to the baseline for the XL model across categories on LegalBench and MMLU.
To analyze the change in performance in more detail, we show the difference to the baseline for the XL model on LegalBench and MMLU across tasks (see Figure 6) and across categories (see Figure 7). We find that FLawN-T5 outperforms baseline Flan-T5 in most LegalBench tasks in most categories. The exception are tasks in the interpretation category, specifically CUAD (Hendrycks et al., 2021c), where the fine-tuned model is actually worse than the baseline by around 10 points on average. A possible explanation could be negative transfer from the instruction tuning data since the task formulations are very different to the instructions in LegalBench. In MAUD (Wang et al., 2023a) and Contract-NLI (Koreeda & Manning, 2021), the instructions are much more similar from LawInstruct to LegalBench, leading to improvements compared to the baseline. On MMLU, most categories and tasks see increases in performance, especially the categories social sciences and other. We find that performance suffers mostly in the STEM category and to some extent in the humanities. Interestingly, the largest drop is in machine learning but the largest rise is in high school computer science. In the humanities, more “hard” disciplines are affected by performance decrease, such as formal logic and logical fallacies.
[…]
4 Ablations
In this section, we present diverse ablations across the starting checkpoints, data mixtures, sampling styles, licences, instruction styles and amount of instruction tuning data during pretraining. We show additional ablations regarding crosslingual transfer from multilingual data in Appendix C.
4.1 Starting Checkpoint
Figure 8: Starting instruction tuning from the Flan-T5 checkpoint improves results across all sizes.
Should you start in-domain instruction tuning from a base model or from an instruction tuned model? $\Rightarrow$ Starting from an instruction tuned model is better across sizes except Small. In Figure 8, we compare instruction tuning from a base T5 and a Flan-T5 model in four different sizes (Small, Base, XL and XXL) (detailed results in Appendix D Table 4). We find that for the larger sizes, the instruction tuned Flan-T5 is a better starting point ($p<0.001$), leading to higher performance on LegalBench. For the Small size the difference is not statistically significant ($p=0.058$). By default, we use the Flan-T5 model as a starting point in all following experiments unless specified otherwise.
4.2 Data Mixture
Figure 9: Accuracy of the Flan-T5 XL model on LegalBench using three data mixtures.
What data mixtures should you choose for in-domain instruction tuning? $\Rightarrow$ Mixing in general instruction tuning datasets is necessary. In Figure 9, we compare instruction tuning with three different data mixtures: lawinstruct, flan2 (Chung et al., 2022), and flan2-lawinstruct (where we sample equally from flan2 and lawinstruct) (detailed results in Appendix D Table 5). Interestingly, when only training on lawinstruct, downstream accuracy drops, possibly due to the instructions in our datasets being formulated differently than the original Flan instructions. Training on flan2 and flan2-lawinstruct leads to an aggregate increase of 7.7 points (48.3 to 56) and 10.8 points (48.3 to 59.1) respectively. By default we use the flan2-lawinstruct mixture in all following experiments unless specified otherwise.
4.3 Sampling Style
Figure 10: Ablation on sampling style and license on English flan2-lawinstruct from the Flan-T5 checkpoint across sizes. Abbreviations: res: licensed for research use (all datasets), comm: commercially friendly licensed, number: sampling by the number of examples per dataset, equal: equally sampling from each dataset
Should we sample each dataset equally or rather by the number of examples? $\Rightarrow$ Sampling by the number of examples generally leads to better performance. In Figure 10, we compare the performance of two sampling styles (equal sampling of each dataset and sampling by the number of examples) across both the research and commercial licensed dataset (detailed results in Appendix D Table 6). For the XL and XXL sizes, sampling by the number of examples is better than equal weight for datasets for both the research and commercial datasets, although not always statistically significant (XL res $p=0.049$, XL comm $p=0.052$, XXL res $p<0.001$, XXL comm $p=0.31$). For the Small size, sampling by the number of examples is better for the research dataset ($p<0.001$) but not for the commercial dataset ($p=0.099$), while there is no difference for the Base size. By default, we sample by the number of examples in all following experiments unless specified otherwise.
4.4 License of Instruction Tuning Datasets
Do we need data licensed non-commercially for good performance? $\Rightarrow$ The commercially licensed data seems to be enough for the larger models. In Figure 10, we compare the performance of two differently licensed datasets (research and commercial licenses) across both sampling each dataset equally and by the number of examples (detailed results in Appendix D Table 6). Except for Small size ($p<0.001$), using more diverse data available only for research shows no significant benefit. By default, we use the commercially licensed dataset in all subsequent experiments unless specified.
4.5 Instruction Style
Are models trained with more diverse instructions better on LegalBench? $\Rightarrow$ Results are mixed, overall just using one instruction is probably sufficient. In Figure 11, we compare the performance of training with just one manually written instruction vs. ten paraphrased instructions with GPT-4 from one seed instruction (detailed results in Appendix D Table 8). For Flan-T5 (see Table 8), for Small, one instruction is better than ten ($p=0.035$); for the other sizes we find no difference. For mT5 (see Figure 11(b)), for Small, one instruction is worse than ten both monolingual ($p=0.005$) and multilingual ($p=0.01$) whereas for XL, ten English instructions underperform one English ($p<0.001$) and ten multilingual ones ($p<0.001$). In aggregate, differences are small without a consistent trend.
(a) English
(b) Multilingual
Figure 11: Ablation on the instruction style on English/multilingual flan2-lawinstruct from the Flan-T5/mT5 checkpoint across all sizes.
4.6 Amount of Instruction Data During Continued Pretraining
How much instruction tuning data should be mixed in during continued pretraining? $\Rightarrow$ Continued pretraining seems to be rather robust with regard to the amount of instruction tuning samples mixed in. In Tables 10, 11, 12 and 13, we investigate the benefit of mixing varying amounts of instruction tuning data in during continued pretraining (detailed results in Appendix D Tables 10, 11, 12 and 13). We compare results on LegalBench of instruction tuning runs after 10K to 90K steps of continued pretraining. For the Small model, the benefit of continued pretraining over just instruction tuning is significant (34.9 for just instruction tuning vs.  40 after continued pretraining). Conversely, for the XL model, continued pretraining often underperforms compared to just instruction tuning. For the XXL model, more instruction tuning samples during continued pretraining improve performance, unlike for the Small and XL models. Across sizes, continued pretraining’s effectiveness appears robust to the number of instruction tuning samples used. [^4] Just mixing instruction tuning data during continued pretraining without more instruction tuning does not improve results.
5 Related Work
Domain-specific pretraining, covering areas such as medicine, law, and science, significantly enhances Language Model performance on related tasks (Beltagy et al., 2019; Gu et al., 2021; Chalkidis et al., 2020). SciBERT (Beltagy et al., 2019), for instance, underwent pretraining on a combination of computer science and biomedical papers, exemplifying this approach in the scientific domain. Other models like PubMedBERT (Gu et al., 2021) and BioBERT (Lee et al., 2020), specifically pretrained on biomedical datasets, have shown improvements in medical NLP tasks (Huang et al., 2019).
5.1 Domain-specific Legal Pretraining
In the legal domain, models such as LegalBERT, pretrained on 12 GB of English legal texts, demonstrated notable success in domain-specific challenges (Chalkidis et al., 2020). CaseLaw-BERT capitalized on the English Harvard Law case corpus spanning from 1965 to 2021 (Zheng et al., 2022), while Niklaus & Giofré (2022) pretrained LongFormer models on the Pile-of-Law (Henderson et al., 2022) using the replaced token detection task (Clark et al., 2020) for enhanced performance. Further advancements were made by Chalkidis et al. (2023), who developed new English legal LMs yielding superior results on LexFiles, a compilation of 11 sub-corpora from six English-speaking legal systems encompassing 19B tokens. Additionally, Niklaus et al. (2023b) introduced a vast multilingual legal corpus, training both monolingual and multilingual legal models to achieve state-of-the-art results on LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023a). Models have also been developed for specific jurisdictions, including the Swiss (Rasiah et al., 2023), Italian (Licari & Comandè, 2022), Romanian (Masala et al., 2021), and Spanish (Gutiérrez-Fandiño et al., 2021) legal systems. Despite the prevalence of smaller encoder-based legal-specific LMs, larger generative models in this space remain scarce. This work seeks to bridge that gap.
5.2 Instruction Tuning
Instruction tuning – the process of finetuning auto-regressive pretrained language models on corpora of reciprocal instruction–response pairs – has emerged as a critical step for building responsive models that are useful for many tasks (Ouyang et al., 2022; Chowdhery et al., 2022; Wei et al., 2022b; Sanh et al., 2022). Some go as far as to claim that this training paradigm is the key to imbuing language models with the generalized capability of zero-shot instruction following behavior (Chung et al., 2022). Instruction tuning refers to few-shot or zero-shot adaptation of large language models to new tasks, where the task is described in natural language in the training examples. Following Wei et al. (2022a), it is common to transform existing datasets into instruction datasets by manually composing templates and filling these with specific examples. It is through these domain-specific training procedures that we build and evaluate legal data adaptation in LLMs.
6 Conclusion and Future Work
We curated LawInstruct, the first instruction tuning dataset for the legal domain by aggregating various high-quality annotated datasets and writing instructions for the different tasks. We used LawInstruct to instruction tune T5 based models, creating FLawN-T5 and a new state-of-the-art on LegalBench in all investigated parameter sizes. We openly release LawInstruct on Hugging Face.
In the future, we would like to extend LawInstruct with more high-quality datasets released after our experiments such as Long-form Legal Question Answering (Louis et al., 2023), Keyphrase Generation (Salaün et al., 2024), Negation Scope Resolution (Christen et al., 2023), or Legal Violation Detection (Bernsohn et al., 2024). Additionally, future work could work with legal professionals to identify and annotate new tasks for the legal domain. Future work could explore the benefits of synthetic data generation in the legal domain, either by generating responses from instructions (Wang et al., 2023c) or reversely, by generating instructions to responses (Köksal et al., 2024), but should take care to do detailed quality checks since hallucinated content may hurt more than improve, especially in the legal domain. Finally, it would be interesting to investigate overlap between the T5 pretraining dataset C4 and the MultiLegalPile to get a better understanding of the potential benefits of continued pretraining.
Acknowledgements
We would like to thank Baq Haidri for help with debugging and John Kirchenbauer, Sang Michael Xie and David Hall for feedback on the project. Thanks to Neel Guha for help with debugging LegalBench and for feedback on analyzing results.
Appendix A Detailed Dataset Description
Figure 12 shows the LawInstruct task type and jurisdiction composition by dataset. Table 2 lists the dataset (and sources), license, language, jurisdiction, task type, subtask, and number of examples for each dataset included in LawInstruct.
Figure 12: Jurisdiction and task type by datasets.
Dataset License Languages Jurisdiction Tasks Subtask Examples Benchmark for Understanding Indian Legal Documents (BUILD) (Kalamkar et al., 2022) Unknown en India Text classification Rhetorical role 28,986 Brazilian Bar Exam (Delfino et al., 2017) Unknown pt Brazil Question answering Bar exam questions 2,130 Brazilian Court Decisions (Lage-Freitas et al., 2022) Unknown pt Brazil Text classification Judgment 3,234 Brazilian Court Decisions (Lage-Freitas et al., 2022) Unknown pt Brazil Text classification Decision Unanimity 1,715 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Multiple choice Judgment 1,225,922 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Judgment 612,961 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Area of law 612,961 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Topic 1,838,883 BVADecisions (Walker et al., 2019) MIT en USA Text classification Rhetorical role 8,818 BVADecisions (Walker et al., 2019) MIT en USA Question answering Relevant rules 2 CAIL 2019 (Xiao et al., 2019) Unknown zh China Question answering Chinese legal case questions 39,333 CAIL 2022 (CAIL, 2022) Unknown zh China Text classification Charge/crime 10,448 CAIL 2022 (CAIL, 2022) Unknown zh China Argument & counter-argument 5,224 CAIL 2022 (CAIL, 2022) Unknown zh China Question answering Response to argument 5,224 Case Briefs (Case briefs, ) CC BY-NC en USA Question answering Legal analysis of facts 2,619 CaseHOLD (Zheng et al., 2021) CC-BY en USA Multiple choice Legal holding statements 45,000 Change My View (Tan et al., 2016) Unknown en N/A Argument & counter-argument 3,456 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Question generation Entailed question 1,774 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Natural language inference Passage entailment 125,954 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Question answering Relevant legal rule 1,774 ContractNLI (Koreeda & Manning, 2021) CC BY-NC en USA Natural language inference Premise hypothesis entailment 14,010 COVID-19 Emergency Measures (EXCEPTIUS) (Tziafas et al., 2021) Unknown en, fr, hu, it, nb, nl, pl EU Text classification Measure type 3,312 European Court of Human Rights (ECtHR) (Chalkidis et al., 2021b) CC BY-NC-SA 4.0 en EU Text classification (multi-label) Violated article 9,000 European Court of Human Rights (ECtHR) (Chalkidis et al., 2021b) CC BY-NC-SA 4.0 en EU Text classification (multi-label) Allegedly violated article 9,000 EOIR (Henderson et al., 2022) CC BY-NC-SA 4.0 en USA Text classification Pseudonymity 8,089 EURLEX (Chalkidis et al., 2019) CC BY-SA 4.0 en EU Text classification EuroVoc core concepts 55,000 EUR-Lex-Sum (Aumiller et al., 2022) CC BY 4.0 24 EU langs EU Summarization EU Legal Acts 22,989 German Argument Mining (Urchs et al., 2021) CC BY 4.0 de Germany Text classification Argumentative function 19,271 German Rental Agreements (Steinberger et al., 2006) Unknown de Germany Text classification Semantic type 3,292 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Volume (coarse thematic topic) 28,536 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Chapter (intermediate thematic topic) 28,536 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Subject (fine-grain thematic topic) 28,536 Greek Legal NER (elNER) (Bartziokas et al., 2020) CC BY-NC-SA 4.0 el Greece Named entity recognition Greek legal entities 17,699 ILDC (Malik et al., 2021) CC BY-NC en India Text classification Judgment 37,387 International Citizenship Law (Vink et al., 2021) CC BY 4.0 en International Question answering Citizenship acquisition 6,460 International Citizenship Law (Vink et al., 2021) CC BY 4.0 en International Question answering Citizenship loss 2,850 JEC-QA (Zhong et al., 2020) CC BY-NC-ND zh China Multiple choice National Judicial Examination of China 21,072 Korean Legal QA (Jeon, 2021) Academic use only ko South Korea Question answering Relevant law 1,830 LawngNLI (Bruno & Roth, 2022) MIT en USA Natural language inference Premise hypothesis entailment 1,142,304 LBOX OPEN (Hwang et al., 2022) CC BY-NC ko South Korea Text classification Judgment 12,142 LBOX OPEN (Hwang et al., 2022) CC BY-NC ko South Korea Text classification Relevant statutes 13,317 LEDGAR (Tuggener et al., 2020) CC BY-NC en USA Text classification Contract provision category 60,000 Legal Case Document Summarization (Shukla et al., 2022; Bhattacharya et al., 2019) CC BY-SA en India Summarization Indian Supreme Court 7,080 Legal Case Summarization (Shukla et al., 2022; Bhattacharya et al., 2019) CC BY-SA en UK Summarization UK Supreme Court 693 LegalNERo (Pais et al., 2021) CC0 1.0 ro Romania Named entity recognition Romanian legal entities 7,552 LegalQA (LegalQA, ) Unknown zh China Question answering Legal advice 21,946 LeNER-Br (Luz de Araujo et al., 2018) Unknown pt Brazil Named entity recognition Brazilian legal entities 7,828 Littleton (Basu et al., 2022) MIT en USA Question answering Relevant future interests 131 Littleton (Basu et al., 2022) MIT en USA Question answering Event graph 143 MAPA (de Gibert Bonet et al., 2022) CC BY-NC 4.0 24 EU langs EU Named entity recognition Coarse-grained 27,823 MAPA (de Gibert Bonet et al., 2022) CC BY-NC 4.0 24 EU langs EU Named entity recognition Fine-grained 27,823 MAUD (Wang et al., 2023a) CC BY en USA Multiple choice Merger agreement questions 10,751 MAUD (Wang et al., 2023a) CC BY en USA Text classification Deal point category 25,827 MAUD (Wang et al., 2023a) CC BY en USA Text classification Question type 25,827 MAUD (Wang et al., 2023a) CC BY en USA Text classification Text type 25,827 Mining Legal Arguments (Habernal et al., 2022) Apache-2.0 en EU Named entity recognition Actors 31,852 Mining Legal Arguments (Habernal et al., 2022) Apache-2.0 en EU Named entity recognition Argument type 31,852 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (coarse level) 1,265,000 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (intermediate level) 911,798 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (fine-grain level) 1,265,000 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Long to short 2,210 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Long to tiny 1,130 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Short to tiny 1,129 Natural Instructions (BillSum) (Kornilova & Eidelman, 2019) CC0 1.0 en USA Summarization U.S Congressional and California state bills 25,200 Natural Instructions (CAIL 2018) (Xiao et al., 2018) Unknown zh China Question answering Judgment 5,988 Natural Instructions (CaseHOLD) (Zheng et al., 2021) CC-BY en USA Multiple choice Correct answer 5,988 Natural Instructions (CaseHOLD) (Zheng et al., 2021) CC-BY en USA Multiple choice Incorrect answer 5,988 Natural Instructions (CUAD) (Hendrycks et al., 2021c) CC BY 4.0 en Question answering Information relevant for contract review 2,442 Natural Instructions (CUAD) (Hendrycks et al., 2021c) CC BY 4.0 en USA Question generation Questions relevant for contract review 2,442 Natural Instructions (EURLEX) (Chalkidis et al., 2019) CC BY-SA 4.0 en EU Text classification Regulation, decisions, or directive 5,850 Natural Instructions (EURLEX) (Aumiller et al., 2022) CC BY-SA 4.0 en EU Summarization EU Legal Acts 3,900 Natural Instructions (OPP-115) (Wilson et al., 2016) CC BY-NC en USA Question answering Type of information used by website 18,480 Natural Instructions (OPP-115) (Wilson et al., 2016) CC BY-NC en USA Question answering Purpose of privacy policy 18,474 Natural Instructions (Overruling) (Zheng et al., 2021) Unknown en USA Text classification Sentence is overruling 14,370 OLC Memos (Henderson et al., 2022) CC BY-NC en USA Question answering Write a legal research memo 1,038 Online ToS (Drawzeski et al., 2021) CC BY-NC 2.5 de, en, it, pt Unknown Text classification Clause topic 19,942 Online ToS (Drawzeski et al., 2021) CC BY-NC 2.5 de, en, it, pt Unknown Text classification Unfair contractual term type 2,074 Plain English Contracts Summarization (Manor & Li, 2019) Unknown en USA Summarization Software licenses, ToS 446 PrivacyQA (Ravichander et al., 2019) MIT en Unknown Question answering Contents of privacy policies 185,200 PrivacySummarization (Keymanesh et al., 2020) MIT en USA Summarization Privacy policies, ToS, and cookie policies 5,751 RedditLegalQA (Henderson et al., 2022) CC BY 4.0 en Unknown Question answering Legal advice from r/legaladvice 192,953 Sara (Holzenberger et al., 2020) Unknown en USA Natural language entailment Fact entailment 176 Sara (Holzenberger et al., 2020) Unknown en USA Question answering Tax liability 160 SaraProlog (Holzenberger et al., 2020) Unknown en USA Question answering Fact pattern to prolog code 376 SaraProlog (Holzenberger et al., 2020) Unknown en USA Question answering Tax statute to prolog code 9 Short Answer Feedback (Filighera et al., 2022) CC BY 4.0 de Germany Question answering Answer question about German law 1,596 Short Answer Feedback (Filighera et al., 2022) CC BY 4.0 de Germany Question answering Feedback rating for answer 1,596 Spanish Labor Law (Calleja et al., 2021) CC BY 4.0 es Spain Extractive question answering Answer question about Spanish labor law 111 StackExchange Questions (Law) (Law Stack Exchange, ) CC BY-SA en Unknown Question answering Online legal forum 10,158 The Supreme Court Database (Spaeth et al., 2020) CC BY-NC 3.0 en USA Text classification Issue areas 5,000 Swiss Federal Supreme Court (Rasiah et al., 2023) CC BY 4.0 de, fr Text generation Case considerations sections (lower court) 26 Swiss Courts (Rasiah et al., 2023) CC BY 4.0 de, fr, it Switzerland Text generation Case considerations sections (same court) 234,313 Swiss Federal Supreme Court (Rasiah et al., 2023) CC BY 4.0 de, fr, it Switzerland Text classification Case criticality (based on citations) 91,075 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Multiple choice Judgment 477,636 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it Switzerland Text classification Judgment 385,719 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Text classification Area of law 18,162 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Text classification Subarea of law 18,162 Swiss Federal Supreme Court (Leading Decisions) (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it Switzerland Text classification Location (canton, region) 42,342 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, fr, it, rm Switzerland Text classification Abbreviation 11,045 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Canton 35,698 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Short description 3,747 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Title 35,359 Thai Supreme Court Cases (TSCC)(Thanh et al., 2021) Academic use only th Thailand Question answering Relevant legal articles (Thai Criminal Code) 2,883 Turkish Constitutional Court (Mumcuoğlu et al., 2021) CC BY 4.0 tr Turkey Multiple choice Judgment 1,804 Turkish Constitutional Court (Mumcuoğlu et al., 2021) CC BY 4.0 tr Turkey Text classification Judgment 902 Unfair ToS (Lippi et al., 2019) Unknown en USA Text classification (multi-label) Unfair contractual term type 5,532 U.S Class Actions (Semo et al., 2022) GPL-3.0 en USA Text classification Judgment 3,000 Valid Wills (Kwak et al., 2022) Unknown en USA Text classification Statement supported by law/condition 1,512
Table 2: Overview of the LawInstruct datasets. The 24 EU langs are bg, cs, da, de, el, en, es, et, fi, fr, ga, hu, it, lt, lv, mt, nl, pt, ro, sv, sk. Abbreviations: Terms of Service (ToS)
Appendix B Detailed Experimental Setup
B.1 Inexplicable Behaviour at the XXL size
We spent considerable effort, including joint debugging with the authors of LegalBench, to reproduce their results. We double checked that the prompts, decoding hyperparameters and general setup are consistent. We conjecture, that the conversion of the Flan-T5 weights as done by Hugging Face on their hub leads to different behavior when running the models with T5X on TPUs (our setup) vs running them with Hugging Face transformers and PyTorch on NVIDIA GPUs (original LegalBench setup) [^5] Similar issues are mentioned in this issue: https://github.com/PiotrNawrot/nanoT5/issues/25.
The XXL mT5 model did not train stably in the continued pretraining phase despite heavy hyperparameter tuning.
B.2 Evaluation
We excluded any legal tasks occurring in MMLU from LawInstruct. However, there is some overlap regarding the tasks included in LawInstruct and in LegalBench because high-quality legal tasks are rare. To control for these overlapping tasks, we evaluate on two versions of LegalBench holding out tasks by the datasets or tasks occurring in LawInstruct respectively.
B.2.1 LegalBench Dataset Held Out
If the source dataset of the LegalBench task occurs in LawInstruct, we remove it from the evaluation. Below, we list which tasks are overlapping. Overall 100 tasks are held out, so 61 tasks are remaining for LegalBench evaluation.
  • ContractNLI
    • LawInstruct: ContractNLI-contract_nli
    • LegalBench: contract_nli_*
  • CUAD
    • LawInstruct: NaturalInstructionsLegal-cuad_answer_generation, NaturalInstructionsLegal-cuad_question_generation
    • LegalBench: cuad_*
  • GLOBALCIT Citizenship Law Dataset
    • LawInstruct: InternationalCitizenshipLawQuestions-international_citizenship_law_questions_mode_acq, InternationalCitizenshipLawQuestions-international_citizenship_law_questions_mode_loss
    • LegalBench: international_citizenship_questions
  • MAUD
    • LawInstruct: MAUD-answer, MAUD-category, MAUD-question, MAUD-text_type
    • LegalBench: maud_*
  • OPP-115 (Online Privacy Policies, set of 115) Corpus
    • LawInstruct: NaturalInstructionsLegal-online_privacy_policy_text_information_type_generation, NaturalInstructionsLegal-online_privacy_policy_text_purpose_answer_generation
    • LegalBench: opp_115_*
  • Overruling
    • LawInstruct: NaturalInstructionsLegal-overruling_legal_classification
    • LegalBench: overruling
  • PrivacyQA
    • LawInstruct: PrivacyQA-privacy_qa
    • LegalBench: privacy_policy_qa
      Note: The LegalBench privacy_policy_entailment Source field is currently incorrectly linked to this dataset (PrivacyQA), but is derived from a different dataset (APP-350 Corpus).
  • StAtutory Reasoning Assessment (SARA)
    • LawInstruct: Sara-sara_entailment, Sara-sara_tax_liability, SaraProlog-sara_prolog_facts, SaraProlog-sara_prolog_statute
    • LegalBench: sara_* (built off of SARA v2)
  • Unfair Terms of Service
    • LawInstruct: This dataset is duplicated between LexGLUE-unfair_tos, LEXTREME-online_terms_of_service_clause_topics (multilingual version), LEXTREME-online_terms_of_service_unfairness_levels (multilingual version)
    • LegalBench: unfair_tos
B.2.2 LegalBench Task Held Out
We additionally catalog instructions which train the LLM for a task captured in LegalBench. It is not necessary that the instruction-response pair in LawInstruct contain data from LegalBench, just that they are about similar legal tasks (e.g., classifying choice-of-forum provisions). Below, we list which tasks are overlapping. Overall 64 tasks are held out, so 97 tasks are remaining for LegalBench evaluation.
  • Rhetorical Role Labeling (Evaluates LLM’s ability to annotate the function of a piece of legal text.)
    • LawInstruct: bva_decisions_label, indian_text_segmentation, german_argument_mining
    • LegalBench: function_of_decision_section, oral_argument_question_purpose
  • Civil Procedure Questions (Evaluates LLM’s application of basic laws of US jurisdiction.)
    • LawInstruct: civipro_questions_generate_*
    • LegalBench: diversity_*, personal_jurisdiction
  • Legal Entailment (Evaluates LLM’s understanding of logical relationships in legal texts.)
    • LawInstruct: coliee_task3_passage_entailment, contract_nli, lawng_nli_entailment
    • LegalBench: contract_nli_*
  • Contractual Clause Classification (Evaluates LLM’s ability to classify contractual clauses.)
    • LawInstruct: unfair_tos, german_rental_agreements
    • LegalBench: cuad_*, jcrew_blocker, unfair_tos, contract_qa
Appendix C Additional Ablations
C.1 Crosslingual Transfer from Multilingual Data
Figure 13: Ablation on the language and license on flan2-lawinstruct from the mT5 checkpoint across all sizes, sampling by the number of examples.
Is there crosslingual transfer from multilingual data? $\Rightarrow$ On the English LegalBench, we do not see any crosslingual transfer. In Figure 13, we compare the performance of the complete multilingual instruction dataset and the English subset across two differently licensed datasets (research and commercial licenses). We see no statistically significant difference between the multilingual training and the English training. We also see no difference between the differently licensed datasets. This means that just training on the commercial subset is enough. We show detailed results on individual LegalBench categories in Appendix D Table 7. Per default we use the English dataset in all following experiments unless specified otherwise.
Appendix D Detailed Results
Table 3: Baseline results on LegalBench.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Flan-T5 XXL (ours) 36.1 18.8 25.2 35.1 41.1 31.3
Flan-T5-XXL (Guha et al., 2023) 66.0 36.0 63.3 64.4 70.7 60.1
LLaMA-2-13B (Guha et al., 2023) 50.2 37.7 59.3 50.9 54.9 50.6
OPT-13B (Guha et al., 2023) 52.9 28.4 45.0 45.1 43.2 42.9
Vicuna-13B-16k (Guha et al., 2023) 34.3 29.4 34.9 40.0 30.1 33.7
WizardLM-13B (Guha et al., 2023) 24.1 38.0 62.6 50.9 59.8 47.1
Flan-T5 XL (ours) 53.5 32.1 46.8 58.7 59.6 50.1
Flan-T5-XL (Guha et al., 2023) 56.8 31.7 52.1 51.4 67.4 51.9
BLOOM-3B (Guha et al., 2023) 47.4 20.6 45.0 45.0 36.4 38.9
Incite-3B-Instruct (Guha et al., 2023) 51.1 26.9 47.4 49.6 40.2 43.0
OPT-2.7B (Guha et al., 2023) 53.7 22.2 46.0 44.4 39.8 41.2
Flan-T5 Base (ours) 44.7 18.0 20.9 28.9 37.0 29.9
Flan-T5 Small (ours) 0.3 30.4 39.8 28.2 27.7 25.3
Table 4: The T5 and Flan-T5 models finetuned on flan2-lawinstruct in four sizes.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Small T5 45.5 ± 13.2 25.0 ± 28.9 25.6 ± 27.4 18.6 ± 23.6 32.9 ± 26.8 29.5 ± 10.3
Small Flan-T5 25.0 ± 22.0 38.1 ± 25.4 33.1 ± 24.4 20.6 ± 26.4 40.7 ± 19.5 31.5 ± 8.5
Base T5 49.8 ± 0.7 38.1 ± 25.4 34.0 ± 23.3 21.3 ± 22.8 38.0 ± 19.4 36.2 ± 10.2
Base Flan-T5 50.3 ± 2.4 38.8 ± 25.9 34.0 ± 22.4 43.0 ± 21.1 54.1 ± 13.0 44.1 ± 8.2
XL T5 47.8 ± 12.5 37.5 ± 25.0 38.2 ± 15.5 28.6 ± 25.1 49.4 ± 8.1 40.3 ± 8.5
XL Flan-T5 65.7 ± 15.2 45.1 ± 30.3 49.0 ± 23.5 56.8 ± 18.8 79.0 ± 11.4 59.1 ± 13.6
XXL T5 52.7 ± 6.8 38.5 ± 25.7 50.0 ± 22.8 44.9 ± 25.2 70.7 ± 20.5 51.4 ± 12.1
XXL Flan-T5 55.2 ± 23.7 46.3 ± 31.6 56.1 ± 29.1 57.7 ± 19.8 84.6 ± 9.6 60.0 ± 14.4
Table 5: The Flan-T5 models finetuned on three different data mixtures.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Small baseline 0.3 ± 0.7 30.4 ± 20.3 23.8 ± 25.0 16.9 ± 21.1 32.8 ± 21.4 20.8 ± 13.0
Small lawinstruct 0.0 ± 0.1 15.9 ± 23.9 10.7 ± 22.7 10.5 ± 19.8 18.6 ± 25.7 11.1 ± 7.1
Small flan2 28.2 ± 22.4 37.8 ± 25.3 35.1 ± 24.2 22.6 ± 23.3 40.5 ± 19.4 32.8 ± 7.3
Small flan2-lawinstruct 25.0 ± 22.0 38.1 ± 25.4 33.1 ± 24.4 20.6 ± 26.4 40.7 ± 19.5 31.5 ± 8.5
Base baseline 44.7 ± 12.4 18.0 ± 23.6 36.0 ± 23.8 15.6 ± 19.9 42.7 ± 19.8 31.4 ± 13.8
Base lawinstruct 14.6 ± 14.7 22.3 ± 26.3 30.2 ± 22.6 19.7 ± 26.0 17.8 ± 27.4 20.9 ± 5.9
Base flan2 47.2 ± 4.3 37.6 ± 25.0 28.6 ± 23.4 32.5 ± 21.9 54.4 ± 16.3 40.0 ± 10.6
Base flan2-lawinstruct 50.3 ± 2.4 38.8 ± 25.9 34.0 ± 22.4 43.0 ± 21.1 54.1 ± 13.0 44.1 ± 8.2
XL baseline 53.5 ± 6.0 32.1 ± 24.6 38.2 ± 22.4 49.8 ± 22.6 68.1 ± 20.1 48.3 ± 14.0
XL lawinstruct 54.5 ± 7.7 30.2 ± 35.1 42.9 ± 20.8 39.8 ± 30.8 63.7 ± 14.1 46.2 ± 13.1
XL flan2 65.5 ± 14.6 40.6 ± 27.7 52.0 ± 25.6 53.0 ± 21.9 74.0 ± 20.8 57.0 ± 13.0
XL flan2-lawinstruct 65.7 ± 15.2 45.1 ± 30.3 49.0 ± 23.5 56.8 ± 18.8 79.0 ± 11.4 59.1 ± 13.6
XXL baseline 36.1 ± 21.5 18.8 ± 24.6 39.4 ± 32.1 25.7 ± 24.2 47.6 ± 14.0 33.5 ± 11.4
XXL lawinstruct 54.1 ± 7.2 37.7 ± 27.2 53.2 ± 32.6 46.7 ± 25.0 73.7 ± 15.1 53.1 ± 13.3
XXL flan2 64.0 ± 12.6 44.7 ± 31.4 56.4 ± 27.7 55.5 ± 20.2 81.3 ± 9.7 60.4 ± 13.6
XXL flan2-lawinstruct 55.2 ± 23.7 46.3 ± 31.6 56.1 ± 29.1 57.7 ± 19.8 84.6 ± 9.6 60.0 ± 14.4
Table 6: Flan-T5 models finetuned on four different licence-sampling style configurations.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Small res-number 50.3 ± 1.3 38.2 ± 25.5 34.9 ± 25.6 21.3 ± 26.6 45.3 ± 22.0 38.0 ± 11.1
Small res-equal 34.9 ± 21.2 37.5 ± 25.0 33.0 ± 25.3 21.1 ± 25.1 43.8 ± 19.2 34.1 ± 8.3
Small comm-number 25.0 ± 22.0 38.1 ± 25.4 33.1 ± 24.4 20.6 ± 26.4 40.7 ± 19.5 31.5 ± 8.5
Small comm-equal 31.6 ± 25.1 37.2 ± 24.8 33.6 ± 22.8 20.2 ± 24.0 42.6 ± 21.3 33.1 ± 8.3
Base res-number 49.8 ± 3.2 38.1 ± 25.4 36.0 ± 23.8 42.8 ± 21.3 49.5 ± 12.1 43.3 ± 6.3
Base res-equal 48.9 ± 3.8 39.4 ± 26.3 38.4 ± 25.6 36.6 ± 19.8 53.4 ± 18.3 43.4 ± 7.4
Base comm-number 50.3 ± 2.4 38.8 ± 25.9 34.0 ± 22.4 43.0 ± 21.1 54.1 ± 13.0 44.1 ± 8.2
Base comm-equal 49.2 ± 2.9 38.5 ± 25.7 36.4 ± 20.3 40.5 ± 19.8 52.6 ± 13.3 43.4 ± 7.1
XL res-number 59.9 ± 10.4 44.2 ± 29.8 53.5 ± 28.0 57.1 ± 20.2 82.4 ± 11.1 59.4 ± 14.2
XL res-equal 58.2 ± 8.4 42.3 ± 28.7 46.6 ± 16.8 55.4 ± 19.3 79.0 ± 11.9 56.3 ± 14.3
XL comm-number 65.7 ± 15.2 45.1 ± 30.3 49.0 ± 23.5 56.8 ± 18.8 79.0 ± 11.4 59.1 ± 13.6
XL comm-equal 59.3 ± 10.4 40.6 ± 27.2 47.7 ± 20.7 54.1 ± 20.0 78.7 ± 11.9 56.1 ± 14.4
XXL res-number 62.9 ± 12.3 46.9 ± 31.7 57.6 ± 30.2 56.7 ± 21.5 82.3 ± 9.3 61.3 ± 13.1
XXL res-equal 54.9 ± 6.3 43.3 ± 30.1 55.5 ± 27.3 55.4 ± 19.2 70.5 ± 11.6 55.9 ± 9.6
XXL comm-number 55.2 ± 23.7 46.3 ± 31.6 56.1 ± 29.1 57.7 ± 19.8 84.6 ± 9.6 60.0 ± 14.4
XXL comm-equal 59.5 ± 13.1 45.7 ± 30.0 54.8 ± 27.6 55.4 ± 19.6 77.2 ± 12.3 58.5 ± 11.6
Table 7: Flan-T5 models finetuned on four different language-license configurations.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Small all-res 46.8 ± 12.2 38.2 ± 24.2 33.8 ± 22.8 20.1 ± 22.0 36.5 ± 21.1 35.1 ± 9.7
Small en-res 50.7 ± 5.9 37.4 ± 24.9 34.0 ± 23.0 18.5 ± 22.9 37.1 ± 23.0 35.5 ± 11.5
Small all-comm 49.7 ± 2.1 38.0 ± 24.1 34.0 ± 23.0 13.2 ± 19.9 35.8 ± 21.8 34.1 ± 13.2
Small en-comm 49.1 ± 13.3 37.5 ± 25.0 34.4 ± 23.3 19.7 ± 23.2 38.2 ± 24.2 35.8 ± 10.6
Base all-res 51.7 ± 4.4 38.7 ± 26.1 33.6 ± 22.7 22.0 ± 23.6 41.8 ± 18.5 37.6 ± 10.9
Base en-res 51.8 ± 5.5 37.5 ± 25.0 37.1 ± 16.5 20.7 ± 22.7 48.0 ± 18.1 39.0 ± 12.1
Base all-comm 51.8 ± 5.2 38.0 ± 25.4 34.3 ± 22.9 23.7 ± 24.7 45.5 ± 12.6 38.7 ± 10.7
Base en-comm 52.0 ± 3.7 37.5 ± 25.0 33.2 ± 22.7 21.9 ± 21.8 46.5 ± 21.2 38.2 ± 11.7
XL all-res 49.9 ± 0.9 37.5 ± 25.0 36.9 ± 18.1 28.3 ± 22.9 52.2 ± 10.7 41.0 ± 9.9
XL en-res 49.9 ± 0.3 37.5 ± 25.0 36.6 ± 18.4 24.8 ± 25.9 50.5 ± 8.6 39.9 ± 10.7
XL all-comm 51.5 ± 2.3 37.5 ± 25.0 36.9 ± 18.1 26.8 ± 24.2 50.7 ± 9.4 40.7 ± 10.4
XL en-comm 49.9 ± 1.0 37.5 ± 25.0 38.3 ± 16.0 27.2 ± 24.3 50.3 ± 9.8 40.6 ± 9.7
XXL all-res 51.5 ± 2.8 38.2 ± 24.2 40.9 ± 18.5 45.3 ± 19.0 56.4 ± 10.4 46.5 ± 7.5
XXL en-res 53.4 ± 5.4 39.0 ± 24.8 40.1 ± 20.5 45.4 ± 20.6 59.0 ± 9.9 47.4 ± 8.7
XXL all-comm 50.6 ± 1.4 38.3 ± 24.3 45.2 ± 22.4 41.0 ± 20.2 58.9 ± 8.7 46.8 ± 8.2
XXL en-comm 52.5 ± 4.1 33.3 ± 27.0 43.9 ± 24.8 47.2 ± 17.8 59.2 ± 16.2 47.2 ± 9.7
Table 8: Flan-T5 models finetuned on two different instruction style configurations.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Small 1-english 28.3 ± 22.1 37.5 ± 25.0 35.3 ± 20.2 21.8 ± 26.5 44.8 ± 17.9 33.6 ± 8.8
Small 10-english 25.0 ± 22.0 38.1 ± 25.4 33.1 ± 24.4 20.6 ± 26.4 40.7 ± 19.5 31.5 ± 8.5
Base 1-english 51.1 ± 6.2 39.0 ± 26.0 36.2 ± 21.6 43.6 ± 21.2 57.6 ± 14.7 45.5 ± 8.8
Base 10-english 50.3 ± 2.4 38.8 ± 25.9 34.0 ± 22.4 43.0 ± 21.1 54.1 ± 13.0 44.1 ± 8.2
XL 1-english 60.6 ± 11.1 42.5 ± 28.8 52.1 ± 24.4 55.0 ± 18.7 81.3 ± 11.1 58.3 ± 14.5
XL 10-english 65.7 ± 15.2 45.1 ± 30.3 49.0 ± 23.5 56.8 ± 18.8 79.0 ± 11.4 59.1 ± 13.6
XXL 1-english 63.0 ± 13.1 43.9 ± 29.7 59.0 ± 30.5 58.1 ± 20.2 80.7 ± 9.9 60.9 ± 13.2
XXL 10-english 55.2 ± 23.7 46.3 ± 31.6 56.1 ± 29.1 57.7 ± 19.8 84.6 ± 9.6 60.0 ± 14.4
Table 9: mT5 models finetuned on three different instruction style configurations.
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Small 1-english 30.2 ± 20.4 39.4 ± 25.1 35.0 ± 24.3 18.3 ± 24.2 37.8 ± 24.4 32.2 ± 8.5
Small 10-english 50.8 ± 3.1 38.4 ± 25.7 33.8 ± 23.6 17.9 ± 23.9 36.0 ± 23.0 35.4 ± 11.8
Small 10-multi 46.5 ± 13.4 39.4 ± 25.1 33.4 ± 23.5 18.2 ± 24.2 36.9 ± 23.5 34.9 ± 10.5
Base 1-english 53.4 ± 5.7 37.5 ± 23.8 34.7 ± 23.7 26.3 ± 23.7 44.3 ± 20.0 39.2 ± 10.2
Base 10-english 52.4 ± 5.1 37.3 ± 23.6 38.0 ± 17.9 21.8 ± 23.0 41.5 ± 20.5 38.2 ± 11.0
Base 10-multi 51.3 ± 3.2 38.0 ± 24.1 34.4 ± 22.7 29.6 ± 21.2 41.7 ± 18.1 39.0 ± 8.2
XL 1-english 51.7 ± 3.4 38.0 ± 24.1 36.9 ± 18.1 36.3 ± 21.7 50.9 ± 8.9 42.7 ± 7.8
XL 10-english 43.6 ± 16.5 38.0 ± 24.1 36.9 ± 18.1 30.9 ± 20.0 45.6 ± 13.8 39.0 ± 5.8
XL 10-multi 51.2 ± 3.3 38.0 ± 24.1 36.9 ± 18.1 31.1 ± 25.4 54.8 ± 12.9 42.4 ± 10.1
Table 10: Flan-T5 Small models with different domain adaptation strategies (amount of IFT data during continued pretraining).
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Baseline 0.3 ± 0.7 30.4 ± 20.3 39.8 ± 20.8 28.2 ± 21.6 27.7 ± 21.9 25.3 ± 13.2
IFT 25.0 ± 22.0 38.1 ± 25.4 43.0 ± 17.1 36.1 ± 26.5 32.6 ± 24.2 34.9 ± 6.0
1-IFT-to-200-PRE+IFT 10K 50.6 ± 4.2 38.2 ± 25.6 44.3 ± 15.6 33.8 ± 23.3 33.7 ± 23.8 40.1 ± 6.5
1-IFT-to-200-PRE+IFT 20K 50.8 ± 2.2 37.9 ± 25.3 44.4 ± 15.7 35.5 ± 25.1 31.9 ± 24.0 40.1 ± 6.7
1-IFT-to-200-PRE+IFT 30K 42.2 ± 16.2 37.3 ± 24.9 39.8 ± 19.4 34.3 ± 23.7 32.4 ± 23.5 37.2 ± 3.6
1-IFT-to-200-PRE+IFT 40K 45.8 ± 10.8 37.7 ± 25.2 39.7 ± 20.8 35.1 ± 24.4 33.4 ± 24.0 38.3 ± 4.3
1-IFT-to-200-PRE+IFT 50K 47.0 ± 8.8 37.4 ± 24.9 38.9 ± 20.7 35.6 ± 24.6 34.1 ± 21.0 38.6 ± 4.5
1-IFT-to-200-PRE+IFT 60K 50.0 ± 0.4 37.1 ± 24.7 39.3 ± 18.7 34.7 ± 23.3 33.8 ± 21.7 39.0 ± 5.8
1-IFT-to-200-PRE+IFT 70K 41.4 ± 16.9 38.4 ± 25.6 38.8 ± 21.1 34.0 ± 22.7 33.8 ± 22.9 37.3 ± 2.9
1-IFT-to-200-PRE+IFT 80K 51.8 ± 3.8 38.2 ± 25.5 38.5 ± 20.9 36.2 ± 22.6 33.4 ± 21.5 39.6 ± 6.4
1-IFT-to-200-PRE+IFT 90K 42.4 ± 16.7 37.9 ± 25.3 39.7 ± 20.3 35.8 ± 23.5 34.1 ± 22.2 38.0 ± 2.9
1-IFT-to-1000-PRE+IFT 10K 42.3 ± 16.1 38.1 ± 25.4 43.9 ± 15.0 33.6 ± 23.8 32.7 ± 24.5 38.1 ± 4.5
1-IFT-to-1000-PRE+IFT 20K 41.7 ± 20.5 37.0 ± 24.7 42.9 ± 16.6 33.1 ± 23.4 33.0 ± 24.6 37.5 ± 4.2
1-IFT-to-1000-PRE+IFT 30K 49.9 ± 0.4 37.8 ± 25.3 40.3 ± 17.7 34.3 ± 24.2 32.4 ± 23.5 38.9 ± 6.1
1-IFT-to-1000-PRE+IFT 40K 51.4 ± 2.7 37.8 ± 25.2 38.9 ± 20.6 34.7 ± 24.4 33.0 ± 22.5 39.2 ± 6.5
1-IFT-to-1000-PRE+IFT 50K 51.6 ± 2.7 37.7 ± 25.2 39.8 ± 18.4 33.7 ± 23.3 33.8 ± 22.4 39.3 ± 6.6
1-IFT-to-1000-PRE+IFT 60K 50.0 ± 0.6 37.5 ± 25.0 40.5 ± 20.2 34.4 ± 23.5 33.2 ± 22.4 39.1 ± 6.0
1-IFT-to-1000-PRE+IFT 70K 50.3 ± 1.4 37.3 ± 24.9 43.1 ± 17.1 34.6 ± 24.6 33.1 ± 22.4 39.7 ± 6.3
1-IFT-to-1000-PRE+IFT 80K 50.6 ± 1.5 37.7 ± 25.2 43.0 ± 17.4 34.0 ± 23.1 32.9 ± 23.0 39.6 ± 6.5
1-IFT-to-1000-PRE+IFT 90K 51.6 ± 2.6 37.0 ± 24.7 40.2 ± 19.2 34.4 ± 24.8 32.9 ± 21.4 39.2 ± 6.7
1-IFT-to-10000-PRE+IFT 10K 46.0 ± 12.1 38.0 ± 25.4 44.4 ± 15.5 33.5 ± 23.3 33.8 ± 24.3 39.1 ± 5.2
1-IFT-to-10000-PRE+IFT 20K 50.5 ± 1.4 37.9 ± 25.3 44.3 ± 15.4 34.9 ± 25.2 32.1 ± 24.0 39.9 ± 6.7
1-IFT-to-10000-PRE+IFT 30K 51.3 ± 4.0 38.2 ± 25.5 40.5 ± 18.1 33.6 ± 23.3 34.7 ± 26.5 39.7 ± 6.3
1-IFT-to-10000-PRE+IFT 40K 52.3 ± 4.4 38.9 ± 26.1 38.8 ± 19.8 33.2 ± 23.0 33.6 ± 25.3 39.4 ± 6.9
1-IFT-to-10000-PRE+IFT 50K 47.3 ± 12.3 37.6 ± 25.1 41.5 ± 17.2 35.1 ± 24.4 32.8 ± 22.2 38.8 ± 5.1
1-IFT-to-10000-PRE+IFT 60K 49.4 ± 2.7 38.1 ± 25.5 39.0 ± 20.6 35.3 ± 24.3 32.2 ± 23.2 38.8 ± 5.8
1-IFT-to-10000-PRE+IFT 70K 49.2 ± 13.9 37.7 ± 25.2 42.1 ± 16.2 33.2 ± 23.1 33.8 ± 24.3 39.2 ± 5.9
1-IFT-to-10000-PRE+IFT 80K 51.4 ± 7.0 37.5 ± 25.0 42.5 ± 16.0 33.5 ± 22.4 32.7 ± 22.4 39.5 ± 6.9
1-IFT-to-10000-PRE+IFT 90K 44.1 ± 20.2 37.5 ± 25.0 43.0 ± 16.4 33.6 ± 22.3 33.0 ± 21.9 38.2 ± 4.6
ONLY-PRE+IFT 10K 51.1 ± 3.1 37.9 ± 25.3 44.9 ± 16.9 33.8 ± 23.6 34.6 ± 24.7 40.5 ± 6.6
ONLY-PRE+IFT 20K 51.4 ± 4.4 38.1 ± 25.5 43.9 ± 14.0 34.1 ± 25.1 33.2 ± 25.3 40.2 ± 6.8
ONLY-PRE+IFT 30K 43.0 ± 17.8 37.9 ± 25.4 42.2 ± 16.2 35.1 ± 25.6 32.4 ± 23.6 38.1 ± 4.1
ONLY-PRE+IFT 40K 47.1 ± 12.5 38.4 ± 25.6 42.5 ± 16.6 34.9 ± 25.0 32.9 ± 24.5 39.2 ± 5.1
ONLY-PRE+IFT 50K 42.0 ± 19.2 37.8 ± 25.2 42.3 ± 17.4 34.8 ± 25.1 32.4 ± 23.3 37.8 ± 3.9
ONLY-PRE+IFT 60K 50.6 ± 2.1 37.9 ± 25.3 43.0 ± 16.0 35.6 ± 25.0 32.6 ± 22.9 39.9 ± 6.3
ONLY-PRE+IFT 70K 48.6 ± 7.0 38.1 ± 25.4 42.6 ± 17.0 34.8 ± 24.3 32.6 ± 24.0 39.4 ± 5.7
ONLY-PRE+IFT 80K 51.2 ± 3.4 37.5 ± 25.0 43.7 ± 17.2 33.2 ± 23.1 34.0 ± 25.7 39.9 ± 6.7
ONLY-PRE+IFT 90K 51.5 ± 3.7 37.5 ± 25.0 40.7 ± 17.5 34.7 ± 21.8 33.7 ± 24.4 39.6 ± 6.4
Table 11: Flan-T5 Base models with different domain adaptation strategies (amount of IFT data during continued pretraining).
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Baseline 44.7 ± 12.4 18.0 ± 23.6 20.9 ± 24.8 28.9 ± 21.2 37.0 ± 21.3 29.9 ± 9.9
IFT 50.3 ± 2.4 38.8 ± 25.9 40.5 ± 15.7 49.5 ± 19.1 45.2 ± 22.0 44.9 ± 4.6
1-IFT-to-200-PRE+IFT 10K 50.5 ± 3.2 37.3 ± 24.9 40.7 ± 16.6 47.7 ± 17.7 49.7 ± 20.8 45.2 ± 5.2
1-IFT-to-200-PRE+IFT 20K 50.4 ± 2.2 37.8 ± 25.2 40.9 ± 14.2 48.4 ± 15.9 46.2 ± 24.7 44.7 ± 4.7
1-IFT-to-200-PRE+IFT 30K 49.9 ± 2.6 37.7 ± 25.2 41.2 ± 14.1 45.3 ± 16.1 48.4 ± 20.0 44.5 ± 4.5
1-IFT-to-200-PRE+IFT 40K 49.4 ± 4.3 37.8 ± 25.2 40.4 ± 15.5 47.8 ± 17.2 49.0 ± 20.9 44.9 ± 4.8
1-IFT-to-200-PRE+IFT 50K 51.2 ± 3.9 37.7 ± 25.2 41.2 ± 12.7 45.0 ± 16.0 49.1 ± 20.1 44.8 ± 4.9
1-IFT-to-200-PRE+IFT 60K 50.1 ± 0.9 37.6 ± 25.1 45.1 ± 13.0 44.2 ± 16.0 45.2 ± 18.9 44.4 ± 4.0
1-IFT-to-200-PRE+IFT 70K 51.1 ± 2.7 37.6 ± 25.0 43.4 ± 13.6 45.1 ± 15.4 46.5 ± 21.0 44.7 ± 4.4
1-IFT-to-200-PRE+IFT 80K 50.4 ± 2.3 37.7 ± 25.2 42.2 ± 15.9 45.2 ± 15.7 44.8 ± 22.7 44.1 ± 4.1
1-IFT-to-200-PRE+IFT 90K 51.4 ± 3.6 37.7 ± 25.2 41.6 ± 14.5 42.9 ± 19.0 43.2 ± 21.6 43.4 ± 4.5
1-IFT-to-1000-PRE+IFT 10K 46.8 ± 4.8 38.5 ± 25.7 43.9 ± 13.7 47.6 ± 16.6 45.9 ± 18.0 44.5 ± 3.2
1-IFT-to-1000-PRE+IFT 20K 50.1 ± 2.0 37.8 ± 25.2 43.2 ± 15.0 46.7 ± 15.9 48.2 ± 24.9 45.2 ± 4.3
1-IFT-to-1000-PRE+IFT 30K 50.8 ± 3.3 38.9 ± 26.0 42.3 ± 15.9 49.9 ± 17.6 50.4 ± 21.4 46.5 ± 4.9
1-IFT-to-1000-PRE+IFT 40K 50.1 ± 0.7 38.4 ± 25.7 45.1 ± 12.0 46.6 ± 16.2 48.0 ± 21.4 45.7 ± 4.0
1-IFT-to-1000-PRE+IFT 50K 51.1 ± 3.0 37.7 ± 25.1 41.9 ± 13.8 48.0 ± 19.3 50.1 ± 20.5 45.8 ± 5.1
1-IFT-to-1000-PRE+IFT 60K 49.9 ± 2.3 37.7 ± 25.1 44.2 ± 15.7 46.1 ± 18.3 49.7 ± 22.1 45.5 ± 4.5
1-IFT-to-1000-PRE+IFT 70K 50.5 ± 1.5 38.5 ± 25.7 44.9 ± 16.8 47.9 ± 15.9 49.8 ± 19.2 46.3 ± 4.4
1-IFT-to-1000-PRE+IFT 80K 50.6 ± 2.5 37.9 ± 25.2 42.4 ± 16.6 48.8 ± 19.2 48.7 ± 22.8 45.7 ± 4.8
1-IFT-to-1000-PRE+IFT 90K 50.8 ± 4.2 37.8 ± 25.2 43.4 ± 15.7 45.9 ± 16.9 47.8 ± 22.0 45.1 ± 4.4
1-IFT-to-10000-PRE+IFT 10K 48.8 ± 4.1 38.1 ± 25.4 43.6 ± 13.4 47.4 ± 16.4 47.7 ± 19.6 45.1 ± 3.9
1-IFT-to-10000-PRE+IFT 20K 50.0 ± 2.9 37.7 ± 25.1 41.5 ± 13.6 47.2 ± 18.4 52.0 ± 20.8 45.7 ± 5.3
1-IFT-to-10000-PRE+IFT 30K 50.5 ± 4.6 38.4 ± 25.6 44.3 ± 14.6 48.4 ± 17.3 51.5 ± 20.7 46.6 ± 4.8
1-IFT-to-10000-PRE+IFT 40K 50.2 ± 2.9 37.7 ± 25.1 42.4 ± 16.4 45.6 ± 16.8 49.2 ± 20.7 45.0 ± 4.6
1-IFT-to-10000-PRE+IFT 50K 50.3 ± 2.0 37.4 ± 24.9 41.8 ± 16.2 45.8 ± 17.7 49.3 ± 21.7 44.9 ± 4.8
1-IFT-to-10000-PRE+IFT 60K 49.6 ± 4.5 37.6 ± 25.1 43.7 ± 17.3 43.1 ± 19.3 48.4 ± 22.0 44.5 ± 4.3
1-IFT-to-10000-PRE+IFT 70K 49.6 ± 2.9 37.7 ± 25.1 46.4 ± 16.0 46.9 ± 18.7 50.5 ± 22.2 46.2 ± 4.5
1-IFT-to-10000-PRE+IFT 80K 49.7 ± 3.0 37.7 ± 25.2 45.1 ± 12.2 41.1 ± 18.4 47.7 ± 23.7 44.2 ± 4.3
1-IFT-to-10000-PRE+IFT 90K 50.0 ± 1.8 37.2 ± 24.8 40.6 ± 14.5 41.8 ± 20.0 45.3 ± 22.3 43.0 ± 4.4
ONLY-PRE+IFT 10K 50.7 ± 2.7 37.2 ± 24.8 42.0 ± 16.3 48.0 ± 18.6 47.6 ± 20.8 45.1 ± 4.9
ONLY-PRE+IFT 20K 50.1 ± 2.6 38.2 ± 25.5 41.1 ± 13.7 45.0 ± 19.7 46.7 ± 25.7 44.2 ± 4.2
ONLY-PRE+IFT 30K 50.7 ± 3.6 38.0 ± 25.3 43.3 ± 15.3 44.6 ± 19.0 48.3 ± 21.6 45.0 ± 4.4
ONLY-PRE+IFT 40K 50.4 ± 3.8 38.4 ± 25.6 41.9 ± 14.5 47.4 ± 17.4 46.8 ± 21.4 45.0 ± 4.3
ONLY-PRE+IFT 50K 50.6 ± 2.5 37.5 ± 25.0 41.1 ± 12.8 44.5 ± 18.6 48.2 ± 21.6 44.4 ± 4.7
ONLY-PRE+IFT 60K 49.6 ± 3.4 37.6 ± 25.1 40.4 ± 15.5 47.2 ± 16.7 46.3 ± 21.0 44.2 ± 4.5
ONLY-PRE+IFT 70K 50.6 ± 1.9 38.4 ± 25.6 41.7 ± 13.2 46.1 ± 18.7 45.5 ± 21.9 44.4 ± 4.2
ONLY-PRE+IFT 80K 51.0 ± 3.1 39.2 ± 26.3 42.2 ± 15.7 46.8 ± 18.0 45.3 ± 21.9 44.9 ± 4.0
ONLY-PRE+IFT 90K 50.5 ± 3.8 37.4 ± 25.0 44.3 ± 14.7 43.2 ± 18.1 44.4 ± 22.5 44.0 ± 4.1
Table 12: Flan-T5 XL models with different domain adaptation strategies (amount of IFT data during continued pretraining).
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Baseline 53.5 ± 6.0 32.1 ± 24.6 46.8 ± 15.6 58.7 ± 21.3 59.6 ± 25.6 50.1 ± 10.1
IFT 65.7 ± 15.2 45.1 ± 30.3 49.5 ± 14.2 61.7 ± 17.1 68.6 ± 24.1 58.1 ± 9.2
1-IFT-to-200-PRE+IFT 10K 56.7 ± 6.9 41.8 ± 28.1 55.2 ± 16.9 62.1 ± 18.6 66.8 ± 23.7 56.5 ± 8.4
1-IFT-to-200-PRE+IFT 20K 63.4 ± 13.8 44.2 ± 29.8 52.5 ± 17.4 58.7 ± 16.8 67.0 ± 23.2 57.2 ± 8.1
1-IFT-to-200-PRE+IFT 30K 58.7 ± 10.3 43.6 ± 29.3 56.3 ± 18.2 60.2 ± 18.4 67.9 ± 24.5 57.3 ± 7.9
1-IFT-to-200-PRE+IFT 40K 58.4 ± 9.7 42.3 ± 28.2 54.3 ± 15.2 61.2 ± 18.8 67.5 ± 23.6 56.7 ± 8.4
1-IFT-to-200-PRE+IFT 50K 61.4 ± 13.3 42.2 ± 28.3 51.8 ± 16.3 59.4 ± 17.9 67.3 ± 23.6 56.4 ± 8.7
1-IFT-to-200-PRE+IFT 60K 57.5 ± 8.7 43.6 ± 29.2 53.5 ± 15.8 60.3 ± 17.5 68.2 ± 23.5 56.6 ± 8.1
1-IFT-to-200-PRE+IFT 70K 58.3 ± 10.2 43.1 ± 28.8 54.3 ± 17.9 58.8 ± 18.2 67.6 ± 22.6 56.4 ± 8.0
1-IFT-to-200-PRE+IFT 80K 58.9 ± 11.0 44.9 ± 30.0 51.1 ± 13.2 59.8 ± 17.3 68.5 ± 23.3 56.6 ± 8.1
1-IFT-to-200-PRE+IFT 90K 55.2 ± 6.9 44.4 ± 30.1 51.7 ± 15.6 57.9 ± 17.0 67.7 ± 24.3 55.4 ± 7.6
1-IFT-to-1000-PRE+IFT 10K 61.3 ± 11.8 41.8 ± 28.0 53.4 ± 16.1 60.9 ± 18.8 67.0 ± 23.1 56.9 ± 8.7
1-IFT-to-1000-PRE+IFT 20K 63.3 ± 13.7 44.3 ± 29.6 52.2 ± 17.4 60.7 ± 17.5 67.3 ± 24.6 57.6 ± 8.3
1-IFT-to-1000-PRE+IFT 30K 58.3 ± 9.8 43.4 ± 29.2 54.4 ± 17.1 61.3 ± 20.4 70.2 ± 25.4 57.5 ± 8.8
1-IFT-to-1000-PRE+IFT 40K 62.5 ± 13.2 45.6 ± 30.6 51.3 ± 17.5 60.1 ± 18.9 68.0 ± 25.6 57.5 ± 8.0
1-IFT-to-1000-PRE+IFT 50K 56.8 ± 7.5 44.7 ± 30.2 51.5 ± 14.5 58.9 ± 16.9 69.7 ± 24.9 56.3 ± 8.3
1-IFT-to-1000-PRE+IFT 60K 54.4 ± 5.3 42.2 ± 28.2 52.7 ± 16.3 59.9 ± 17.8 67.1 ± 23.5 55.2 ± 8.2
1-IFT-to-1000-PRE+IFT 70K 59.7 ± 10.8 44.1 ± 29.5 54.5 ± 17.3 59.4 ± 17.6 67.4 ± 23.4 57.0 ± 7.7
1-IFT-to-1000-PRE+IFT 80K 59.8 ± 11.2 41.6 ± 27.9 52.8 ± 17.2 63.5 ± 19.8 67.3 ± 24.5 57.0 ± 9.0
1-IFT-to-1000-PRE+IFT 90K 60.3 ± 10.6 44.3 ± 29.7 50.5 ± 15.4 57.3 ± 15.9 67.3 ± 23.1 55.9 ± 8.0
1-IFT-to-10000-PRE+IFT 10K 60.0 ± 10.2 42.3 ± 28.4 52.7 ± 16.0 61.6 ± 18.3 68.0 ± 22.8 56.9 ± 8.8
1-IFT-to-10000-PRE+IFT 20K 59.5 ± 11.0 42.6 ± 28.5 52.5 ± 15.7 61.6 ± 18.0 68.1 ± 25.0 56.9 ± 8.7
1-IFT-to-10000-PRE+IFT 30K 62.2 ± 12.2 42.3 ± 28.5 53.6 ± 16.7 62.5 ± 20.1 69.2 ± 25.2 57.9 ± 9.3
1-IFT-to-10000-PRE+IFT 40K 59.7 ± 10.1 43.6 ± 29.2 53.1 ± 15.9 62.6 ± 18.9 67.6 ± 23.1 57.3 ± 8.3
1-IFT-to-10000-PRE+IFT 50K 58.8 ± 8.9 42.9 ± 29.1 52.5 ± 16.9 61.1 ± 17.9 64.6 ± 25.0 56.0 ± 7.6
1-IFT-to-10000-PRE+IFT 60K 55.3 ± 5.6 42.1 ± 28.3 52.1 ± 16.6 59.1 ± 19.0 66.4 ± 23.1 55.0 ± 8.0
1-IFT-to-10000-PRE+IFT 70K 60.3 ± 10.0 43.6 ± 29.5 51.8 ± 16.8 61.2 ± 18.5 69.0 ± 24.7 57.2 ± 8.7
1-IFT-to-10000-PRE+IFT 80K 64.7 ± 13.9 44.4 ± 29.9 50.8 ± 16.9 58.4 ± 17.1 70.4 ± 25.8 57.8 ± 9.3
1-IFT-to-10000-PRE+IFT 90K 63.3 ± 13.3 44.8 ± 30.2 51.9 ± 16.3 58.7 ± 16.6 68.2 ± 25.1 57.4 ± 8.3
ONLY-PRE+IFT 10K 62.8 ± 13.6 44.3 ± 29.8 52.0 ± 16.7 58.9 ± 16.2 68.2 ± 23.9 57.2 ± 8.3
ONLY-PRE+IFT 20K 64.0 ± 13.9 42.6 ± 28.7 52.8 ± 15.6 62.0 ± 18.0 68.7 ± 25.0 58.0 ± 9.3
ONLY-PRE+IFT 30K 52.9 ± 15.5 42.0 ± 28.3 51.5 ± 16.0 62.0 ± 18.7 67.3 ± 24.9 55.1 ± 8.8
ONLY-PRE+IFT 40K 60.4 ± 12.2 43.1 ± 29.1 52.4 ± 16.9 60.6 ± 17.5 68.9 ± 23.4 57.1 ± 8.7
ONLY-PRE+IFT 50K 57.4 ± 8.5 42.6 ± 28.8 51.6 ± 15.3 61.2 ± 18.1 70.0 ± 23.8 56.5 ± 9.2
ONLY-PRE+IFT 60K 56.7 ± 7.6 42.5 ± 28.4 52.0 ± 16.3 61.2 ± 17.9 68.8 ± 23.8 56.2 ± 8.8
ONLY-PRE+IFT 70K 57.2 ± 8.5 42.1 ± 28.4 51.5 ± 17.0 60.8 ± 18.1 70.2 ± 24.8 56.3 ± 9.4
ONLY-PRE+IFT 80K 60.3 ± 11.1 42.4 ± 28.4 54.6 ± 16.4 65.1 ± 20.9 69.2 ± 24.8 58.3 ± 9.3
ONLY-PRE+IFT 90K 60.3 ± 12.0 44.4 ± 29.8 52.3 ± 17.1 59.8 ± 17.8 67.8 ± 24.4 56.9 ± 7.9
Table 13: Flan-T5 XXL models with different domain adaptation strategies (amount of IFT data during continued pretraining).
LLM Issue Rule Conclusion Interpretation Rhetorical LegalBench
Baseline 36.1 ± 21.5 18.8 ± 24.6 25.2 ± 26.0 35.1 ± 22.2 41.1 ± 18.4 31.3 ± 8.1
IFT 55.2 ± 23.7 46.3 ± 31.6 56.2 ± 18.3 66.3 ± 19.7 73.8 ± 24.4 59.6 ± 9.5
1-IFT-to-200-PRE+IFT 10K 53.4 ± 16.2 47.9 ± 32.1 58.1 ± 19.5 63.8 ± 17.6 74.2 ± 27.1 59.5 ± 9.0
1-IFT-to-200-PRE+IFT 20K 53.6 ± 3.7 48.9 ± 32.9 58.8 ± 18.7 65.3 ± 17.5 72.0 ± 25.5 59.7 ± 8.2
1-IFT-to-200-PRE+IFT 30K 56.5 ± 18.3 48.9 ± 31.5 60.5 ± 19.9 65.2 ± 18.3 69.5 ± 24.2 60.1 ± 7.1
1-IFT-to-200-PRE+IFT 40K 58.3 ± 20.2 47.3 ± 30.8 57.9 ± 19.1 65.6 ± 18.2 71.3 ± 24.1 60.1 ± 8.1
1-IFT-to-200-PRE+IFT 50K 60.3 ± 12.6 48.4 ± 31.4 63.2 ± 20.2 67.9 ± 18.9 71.4 ± 26.1 62.2 ± 7.9
1-IFT-to-200-PRE+IFT 60K 58.6 ± 20.5 48.5 ± 31.5 60.9 ± 20.7 67.5 ± 19.9 71.0 ± 24.7 61.3 ± 7.8
1-IFT-to-200-PRE+IFT 70K 58.6 ± 10.5 48.5 ± 31.4 60.6 ± 20.4 65.3 ± 18.4 69.3 ± 23.4 60.5 ± 7.0
1-IFT-to-200-PRE+IFT 80K 53.7 ± 16.4 47.8 ± 30.8 58.8 ± 18.2 63.7 ± 17.7 71.3 ± 25.7 59.1 ± 8.1
1-IFT-to-200-PRE+IFT 90K 52.0 ± 14.5 48.8 ± 31.7 59.4 ± 19.6 64.4 ± 17.9 72.3 ± 25.1 59.4 ± 8.5
1-IFT-to-1000-PRE+IFT 10K 41.1 ± 24.2 45.9 ± 30.3 58.2 ± 18.4 65.5 ± 20.2 68.8 ± 25.2 55.9 ± 10.8
1-IFT-to-1000-PRE+IFT 20K 47.7 ± 24.8 48.0 ± 31.1 60.3 ± 20.3 67.2 ± 19.7 70.3 ± 23.8 58.7 ± 9.4
1-IFT-to-1000-PRE+IFT 30K 40.3 ± 28.4 45.5 ± 29.6 62.3 ± 21.1 67.8 ± 21.1 69.3 ± 22.6 57.0 ± 11.9
1-IFT-to-1000-PRE+IFT 40K 44.2 ± 27.4 46.7 ± 29.9 61.9 ± 21.9 68.6 ± 20.7 71.2 ± 24.9 58.5 ± 11.1
1-IFT-to-1000-PRE+IFT 50K 49.7 ± 25.2 49.1 ± 33.1 55.5 ± 19.2 68.2 ± 19.8 71.4 ± 24.3 58.8 ± 9.3
1-IFT-to-1000-PRE+IFT 60K 44.9 ± 22.0 47.6 ± 30.7 57.9 ± 19.4 69.7 ± 21.1 72.1 ± 26.0 58.5 ± 11.1
1-IFT-to-1000-PRE+IFT 70K 40.6 ± 25.0 48.1 ± 31.2 60.5 ± 20.0 68.2 ± 20.5 72.5 ± 24.4 58.0 ± 12.0
1-IFT-to-1000-PRE+IFT 80K 53.8 ± 23.7 47.9 ± 32.4 53.5 ± 17.5 67.1 ± 19.3 71.8 ± 25.9 58.8 ± 9.1
1-IFT-to-1000-PRE+IFT 90K 47.6 ± 23.5 47.1 ± 30.5 60.1 ± 18.9 65.1 ± 24.3 70.3 ± 23.5 58.0 ± 9.3
1-IFT-to-10000-PRE+IFT 10K 49.8 ± 13.6 46.6 ± 30.0 59.0 ± 16.6 64.6 ± 19.3 72.6 ± 24.7 58.5 ± 9.5
1-IFT-to-10000-PRE+IFT 20K 45.2 ± 27.4 46.3 ± 31.2 58.8 ± 20.1 68.1 ± 19.0 71.7 ± 24.1 58.0 ± 10.9
1-IFT-to-10000-PRE+IFT 30K 46.8 ± 24.6 46.0 ± 29.6 62.6 ± 18.4 66.1 ± 18.1 72.1 ± 25.3 58.7 ± 10.5
1-IFT-to-10000-PRE+IFT 40K 56.8 ± 24.5 46.9 ± 30.4 59.1 ± 19.3 68.3 ± 21.1 72.2 ± 26.2 60.7 ± 8.9
1-IFT-to-10000-PRE+IFT 50K 54.5 ± 28.7 43.1 ± 28.1 62.2 ± 19.8 64.2 ± 19.1 70.2 ± 24.3 58.8 ± 9.3
1-IFT-to-10000-PRE+IFT 60K 52.0 ± 16.0 42.0 ± 28.7 60.3 ± 17.4 65.7 ± 19.6 71.3 ± 24.7 58.2 ± 10.3
1-IFT-to-10000-PRE+IFT 70K 52.2 ± 14.7 47.4 ± 30.8 59.2 ± 18.3 66.6 ± 18.5 70.0 ± 24.1 59.1 ± 8.5
1-IFT-to-10000-PRE+IFT 80K 56.5 ± 18.5 44.9 ± 28.9 59.7 ± 17.2 65.3 ± 17.7 72.3 ± 25.6 59.7 ± 9.1
1-IFT-to-10000-PRE+IFT 90K 45.0 ± 17.4 41.5 ± 27.3 56.3 ± 16.3 66.3 ± 18.5 72.1 ± 25.7 56.2 ± 11.8
ONLY-PRE+IFT 10K 49.2 ± 24.4 47.1 ± 30.4 62.0 ± 20.3 66.9 ± 20.4 71.7 ± 25.1 59.4 ± 9.7
ONLY-PRE+IFT 20K 35.6 ± 24.0 46.2 ± 30.0 56.3 ± 17.9 62.3 ± 18.4 68.6 ± 24.2 53.8 ± 11.7
ONLY-PRE+IFT 30K 46.3 ± 28.4 45.7 ± 29.3 56.1 ± 18.5 67.7 ± 19.9 72.1 ± 25.6 57.6 ± 10.8
ONLY-PRE+IFT 40K 48.8 ± 30.3 45.7 ± 29.5 56.6 ± 18.0 68.1 ± 20.0 71.6 ± 26.3 58.1 ± 10.2
ONLY-PRE+IFT 50K 47.5 ± 24.9 47.1 ± 30.2 53.5 ± 16.2 67.1 ± 19.5 71.8 ± 25.4 57.4 ± 10.2
ONLY-PRE+IFT 60K 33.2 ± 23.3 47.8 ± 30.7 55.0 ± 17.9 63.1 ± 19.7 69.3 ± 25.0 53.7 ± 12.6
ONLY-PRE+IFT 70K 42.7 ± 25.9 47.2 ± 30.5 55.9 ± 19.4 60.7 ± 17.5 68.0 ± 23.8 54.9 ± 9.1
ONLY-PRE+IFT 80K 43.7 ± 25.8 46.3 ± 29.9 55.8 ± 17.1 64.8 ± 18.7 71.8 ± 25.9 56.5 ± 10.7
ONLY-PRE+IFT 90K 55.3 ± 16.9 45.2 ± 28.9 60.0 ± 17.0 64.9 ± 20.0 69.0 ± 24.3 58.9 ± 8.2
(a) Dataset Overlap
(b) Task Overlap
Figure 14: Difference to the baseline for the XL model across categories on LegalBench with dataset and task overlap held out respectively.